OpenStudy (anonymous):

|3a+5|=|4a-7| The solution is

OpenStudy (anonymous):

at the absolute we put the bracket once +ve & other -ve ....like that 3a+5=4a-7......So...... a=12 .....or -3a-5=-4a+7.....So..... a=12